OpenStudy (anonymous):

s(x,y)= sqrt(u(x,y)^2 + v(x,y)^2. Use the chain rule to find ∂s/∂x and ∂s/∂y. Problem : u(x,y) = x(1-x)(1-2y) and V(x,y) = y(y-1)(1-2x);

OpenStudy (amistre64):

where you getting into trouble with this?

OpenStudy (anonymous):

like how do you use the given information to solve the problem

OpenStudy (amistre64):

i like to start by reading it, then I apply it ...

OpenStudy (amistre64):

\[s(x,y)= (u(x,y)^2 + v(x,y)^2)^{1/2}\] this is the problem right?

OpenStudy (anonymous):

i have no clue what to do

OpenStudy (apoorvk):

while calculating del(s)/del(y) (NOT same as ds/dy, 'del' means partial differentiation) you just differentiate with respect to y keeping all other stuff (including any other variable like x constant). ditto for del(s)/(del(x)

OpenStudy (anonymous):

no thats the given information

OpenStudy (anonymous):

thats's the dimensions given

OpenStudy (amistre64):

\[sx= (u(x,y)^2 + v(x,y)^2)'^{1/2}*(u(x,y)^2 + v(x,y)^2)'\] \[(u(x,y)^2 + v(x,y)^2)'=u(x,y)'^2 + v(x,y)'^2\] \[u(x,y)'^2=2(u(x,y))*ux\] \[v(x,y)'^2=2(v(x,y))*vx\]

OpenStudy (amistre64):

thats the basic setup for tackling this problem; the rest is just in making it specific

OpenStudy (anonymous):

Problem : u(x,y) = x(1-x)(1-2y) and V(x,y) = y(y-1)(1-2x); what is this then?

OpenStudy (amistre64):

\[s_x=\frac{2u(x,y)u_x + 2v(x,y)v_x}{2(u(x,y)^2 + v(x,y)^2)^{1/2}}\] \[s_x=\frac{u(x,y)u_x + v(x,y)v_x}{(u(x,y)^2 + v(x,y)^2)^{1/2}}\]

OpenStudy (amistre64):

expand out u and v and take the partials like normal

OpenStudy (amistre64):

or play with them as is and product rul them into submission

OpenStudy (anonymous):

thats's tough lol

OpenStudy (amistre64):

it only tough if you let everyone else do it for you ;) give it a shot so I can see where youre having troubles

OpenStudy (amistre64):

\[u_x = x'(1-x)(1-2y)+x(1-x)'(1-2y)+x(1-x)(1-2y)'\] \[u_x = (1-x)(1-2y)+x(-1)(1-2y)+x(1-x)(0)\] not to bad

OpenStudy (anonymous):

wait what did you do here ^

OpenStudy (amistre64):

its a little thing that we learned back in calc1 called the product rule ....

OpenStudy (anonymous):

ohhh

OpenStudy (amistre64):

\[D[fgh]=f'gh+fg'h+fgh'\]

OpenStudy (amistre64):

\[u_y = x'(1-x)(1-2y)+x(1-x)'(1-2y)+x(1-x)(1-2y)'\] \[u_y = 0(1-x)(1-2y)+x(0)(1-2y)+x(1-x)(-2)\] thats the playground for Uy

OpenStudy (amistre64):

so I dont think this issue is so much of not knowing WHAT to do, but rather there is just alot of clutter to figure out WHEN to do it ....

OpenStudy (anonymous):

yea but i still don't get how its gonna solve the problem

OpenStudy (anonymous):

like i am confused

OpenStudy (amistre64):

after youve determined where the pices go, and how they are derived; you just fit them all together

OpenStudy (amistre64):

put ux where ux goes, vx where vx goes; u where u goes, v where v goes; and call it a day :)

OpenStudy (amistre64):

\[\large s_x=\frac{u(x,y)u_x + v(x,y)v_x}{(u(x,y)^2 + v(x,y)^2)^{1/2}}\]

OpenStudy (anonymous):

ohhh ok i get it

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

yep, good luck ;)

OpenStudy (anonymous):

wait i have one more insane problem. this problem is scary

OpenStudy (amistre64):

to clean that one up id prolly shorthand it to: \[\large s_x=\frac{u(u_x) + v(v_x)}{\sqrt{u^2 + v^2}}\] and sy is the same basic set up \[\large s_y=\frac{u(u_y) + v(v_y)}{\sqrt{u^2 + v^2}}\]

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

type up your insane problem to the left and see who bites ;)

OpenStudy (anonymous):

i typed it lol

OpenStudy (anonymous):

can you give the answer for this one

OpenStudy (amistre64):

http://openstudy.com/code-of-conduct "Don't devalue the question/answer process!" >Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem. If there is someplace that you are still struggling with in the problem, please feel free to be specific as to what it is that is not quite clicking.

OpenStudy (anonymous):

lol i think i got it